OpenStudy (anonymous):

Jean and Mark are going to fill a pool with 2 different sized hoses. Jean can fill the pool in 8 hours, while Mark can complete it in 12 hours. Their supervisor thinks that the job will take 10 hours to complete if they work together. Explain each step in solving this equation and determine if the supervisor is correct or not.

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

the supervisor is incorrect right

OpenStudy (loser66):

justify your thought, please, show me your work

OpenStudy (anonymous):

well the equation is \[\frac{ 1 }{ 8} + \frac{ 1 }{ 12} = \frac{ 1 }{ }\]

OpenStudy (anonymous):

\[\frac{ 1 }{ x}\]

OpenStudy (anonymous):

for the last one

OpenStudy (loser66):

next?

OpenStudy (anonymous):

thats where i get stuck

OpenStudy (anonymous):

i don't know what to do from there

OpenStudy (loser66):

to this kind of problem you have to put everything in neat let V is the capacity of the pool so, Jean, in 1 hour , can fill \(\dfrac{1}{8}V\) ok? Mark, in 1 hour , can fill \(\dfrac{1}{12}V\) right?

OpenStudy (anonymous):

yes

OpenStudy (loser66):

If they work together, the volume of water they can fill in the pool in 1 hour is Jean's work + Mark's work = \(\dfrac{1}{8}V +\dfrac{1}{12}V= \dfrac{5}{24}V\) right?

OpenStudy (anonymous):

yea

OpenStudy (loser66):